OpenStudy (anonymous):

What is the 9th term of the geometric sequence 4, 12, 36, ? This is what I have so far: tn = t1 x r ^(n-1) t1 (1st term) = 9 n = # of terms so n = 4 9=4(1) x ratio How would I find the ratio?

OpenStudy (anonymous):

all you do is mulitply each number by 3 4*3=12*3=36

OpenStudy (anonymous):

multiply*

OpenStudy (anonymous):

that's not what the formula says

hartnn (hartnn):

4, 12, 36 whats the 1st term you see ??? its not 9.

OpenStudy (anonymous):

yeah your formula is wrong

OpenStudy (anonymous):

How is 9 not the first term? What is?

OpenStudy (anonymous):

oh.... 41???

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4*?

OpenStudy (anonymous):

your first term is always the first of the sequence that is ... 4

OpenStudy (anonymous):

you are trying to find the 9th term

hartnn (hartnn):

you want 9th term, so n=9 got this ?

OpenStudy (anonymous):

That's what I used!!!

OpenStudy (anonymous):

wait wait wait

OpenStudy (anonymous):

ok

hartnn (hartnn):

now common ratio , r how do you find it ?

OpenStudy (anonymous):

everyone's saying different things so it's rather confusing

OpenStudy (anonymous):

a2/a1? to find the ratio?

hartnn (hartnn):

yes, so, r=... ?

OpenStudy (anonymous):

common ratio , d = 2nd term divide by 1st term

OpenStudy (anonymous):

r*

OpenStudy (anonymous):

a1 = 4?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

12/4?

OpenStudy (anonymous):

= 3

hartnn (hartnn):

yes, r=3 is correct.

OpenStudy (anonymous):

really? do i do that for every problem?

hartnn (hartnn):

yes, tn = t1 x r ^(n-1) t1 (1st term) = 4 n = # of terms so n = 9 r= 3

OpenStudy (anonymous):

ok perfect thanks:) i'll plug in the numbers now

OpenStudy (anonymous):

Just Remember always : For a geometric progression . The ratio between any two successive terms is called the common ratio. The first term a1 is the first of the sequence

OpenStudy (anonymous):

429981696

hartnn (hartnn):

nooo...... how ?

OpenStudy (anonymous):

12^(9-1)

OpenStudy (anonymous):

4(1) x 3 = 12

hartnn (hartnn):

nah... you don't do 4*3 first

hartnn (hartnn):

4 * (3^8 ) first do 3^8 then multiply by 4

OpenStudy (anonymous):